# Confining quantum field theories

### v1

@@ 4 Quantum field theories, old and new
The relation of these two properties, which is, actually, a mathematical, logical equivalence (uniqueness of the vacuum $\Leftrightarrow$ cluster decomposition property) was observed as early as in [7], it is noted in [4], and is also mentioned in [5, 6] in connection with results from operator algebras, such as the GNS construction. All these, as well as subsequent works, however, and eventually quantum field theory textbooks, adopt them as a postulate (a mathematical axiom) in order to build and justify the $S$-matrix construction, which was used to describe the experimental results investigating particle physics. In fact, the early related theorems [7] were derived even before the proposal of the quark model, which was later included in the Standard Model, and its eventual establishment. The weak interactions also, after the spontaneous symmetry breaking mechanism, are expected to have a unique vacuum state, their excitations become massive with vacuum expectation values of Higgs fields, and the related interactions diminish as Yukawa terms, also giving the possibility of an $S$-matrix construction.
The strong interaction, however, with a linearly rising interaction, and confinement, cannot possibly be expected to satisfy the cluster decomposition property, and thus lead to a similar construction. There are no asymptotic states, and there is no such thing as an $S$-matrix. Accordingly, the vacuum state is not unique, there is a non-trivial operator, such that the second (confining) vacuum is an eigenstate of this operator, leading to confinement properties like (8), instead of (9). This is a consequence of the non-linear, self-interaction of the massless, spin-1, gauge bosons.
Although many features of the strong interaction can be seen and investigated with the methods presented here and in previous works, it is stressed that many properties of ordinary quantum field theory do not hold in such theories. Along with the breakdown of the cluster decomposition principle, one has the absence of an $S$-matrix formalism, and even the absence of a globally defined Lagrangian description. Physics is described locally, in terms of local operators in different compact regions of space-time, and can be matched and patched together, for example, in the spirit of a Wilsonian renormalization group, but globally, one does not have the luxury of an $S$-matrix or Lagrangian.
Once we dismiss the Lagrangian formalism as fundamental, we can investigate both new and conventional quantum field theories also in a new light (and also justify and revisit some older interpretations). The physical quantities that are measured are the probabilities that two or more states in a given volume interact; this defines a coupling constant together with a corresponding interaction operator, again in the spirit of the renormalization group, and the conditions under which these operators are patched together in a global Lagrangian can be investigated; singularities, blow-ups, and Landau poles, then may not be considered as problems of the theory per se, but of the Lagrangian construction (much like the infinities of early quantum field theory were remedied with the process of renormalization). Starting by postulating a Lagrangian may be useful and impressive in expressing, predicting, and justifying symmetries of the theory, but is like starting from the conclusion; the problem is how, if, and under what limits and conditions, this formalism can be reached.
On a mathematical sidenote, observables are elements of operator algebras, $A$, states are specifically defined (normalized, etc) elements of the dual space, $A^{*}$, and interactions are (also appropriately and consistently defined) elements of the double dual, $A^{**}$. The natural embedding of a space into its double dual, gives the possibility of defining interactions also as operators (and even there may be leeway; these spaces are not always reflexive). The various “coupling constants” then, are merely normalizations and initial values of elements of $A^{**}$ and, of course, are not constant, as is already well-known in quantum physics. These mathematical operations of taking the dual spaces, and the associated weak topologies involved, may seem “high-brow” or esoteric, but are, actually, the way that the physical measurements and operations are performed [6].
The weak interaction is also non-Abelian, but after spontaneous symmetry breaking, the Coleman-Weinberg potential that is generated by the mechanism of [1] does not admit a second, non-zero extremum, and does not change the unique vacuum structure, although a full treatment is also necessary.
It is also well-known, that the $S$-matrix in these conventional theories with a unique vacuum, usually starts with a tree contribution that is essentially a classical scattering result, and the role of relativistic, quantum physics is mainly to calculate higher order corrections in this result. In new, non-trivial, confining theories with multiple vacua, the phenomena are intrinsically quantum, and new experimental, as well as theoretical methods need to be devised.
It should also be stressed that, although the quantum electromagnetic and weak interactions are expected to admit a unique vacuum and a well-defined $S$-matrix, any other theory that includes the strong interaction, unified or not, including the Standard Model and its extensions, does not admit a conventional interpretation, except maybe in some approximations, that also need to be specified. An interesting discussion of relative scales and confining theories at various limits appears in [8], although in a different context.
It seems very likely, that the fourth physical interaction, gravity, also highly non-linear, would admit a similar, non-trivial structure, with multiple and mixed vacuum states, and confinement properties related to cosmological and black hole horizons, inflation, information puzzles, even a quantum version of Mach’s principle. One can start in a similar manner and consider the constraints of the gravitational interaction, utilize auxiliary fields in order to express them and investigate the generation and stability of new vacua. A more daring possibility, that is perhaps worth some investigation, would start with the realization that the absence of a conventional quantum field theory formalism and an $S$-matrix, also nullifies the traditional theorems that utilize its properties and symmetries (some versions of them may hold, of course, but need to be re-examined). One may then attempt to relate, couple, or somehow connect, an intrinsic symmetry (like the non-Abelian, weak interaction, or a new such interaction) to a subgroup of the Lorentz symmetry group of the spin-2 field. Instead of a Higgs mechanism of mass generation via spontaneous symmetry breaking with a vacuum expectation value, then one would have a quantum generation of a mass scale as an eigenvalue of a new vacuum state and, as a bonus, include gravity in the process.
As is usually stated in a conclusion, these physical and mathematical investigations will be considered in future works.

### v2

@@ 4 Confining and asymptotic quantum field theories
The relation of these two properties, which is, actually, a mathematical, logical equivalence (uniqueness of the vacuum $\Leftrightarrow$ cluster decomposition property) was observed as early as in [7], it is noted in [4], and is also mentioned in [5, 6] in connection with results from operator algebras, such as the GNS construction. All these, as well as subsequent works, however, and eventually quantum field theory textbooks, adopt them as a postulate (a mathematical axiom) in order to build and justify the $S$-matrix construction, which was used to describe the experimental results investigating particle physics. In fact, the early related theorems [7] were derived even before the proposal of the quark model, which was later included in the Standard Model, and its eventual establishment.
The strong interaction, however, with a linearly rising interaction, and confinement, cannot possibly be expected to satisfy the cluster decomposition property, and thus lead to a similar construction. There are no asymptotic states, and there is no such thing as an $S$-matrix involving elementary gluons and quarks at distances larger than the scale of the theory. An interesting and useful discussion of relative scales and confining theories at various limits appears in [8], although it was in a different context, and should be interpreted now with the condition of the mixed vacuum state.
In the strong interaction described by Yang-Mills theory, the vacuum is a mixed state, with the asymptotic and confining vacua $\Omega_{0}$ and $\Omega_{\mu}$, and the Hilbert space is a direct sum ${\cal H}_{0}\oplus{\cal H}_{\mu}$, so that the Hamiltonian of the theory is also decomposed as
$\begin{pmatrix}H_{0}&0\\ 0&H_{\mu}\end{pmatrix}$ (15)
The existence and stability of the solitons connecting the two vacua enlarges the Hilbert space to ${\cal H}_{0}\oplus{\cal H}_{\mu}\oplus{\cal H}_{\rm sol}$, and the interactions can be schematically described by the Hamiltonian
$\begin{pmatrix}H_{0}&0&H_{\rm s0}\\ 0&H_{\mu}&H_{\rm s1}\\ H_{\rm s0}&H_{\rm s1}&H_{\rm sol}\end{pmatrix}$ (16)
From the previous analysis, only $H_{0}$ and $H_{\mu}$ have been completely determined, as in (3) in the work of [2], with the appropriate commutation relations and Feynman rules. The interactions involving solitons, $H_{\rm sol}$, and the off-diagonal terms, can be calculated in principle, but the analytical treatment is limited to large soliton separations.
It should be stressed that the interactions are unitary and Lorentz invariant. Because of confinement, however, as expressed in general in (14), and as we know from everyday practice, the manifestation and verification of causality and unitarity are performed in mixed and composite states of baryons, protons, etc., instead of elementary quark and gluon states.
We have a unitary evolution, as described schematically in (15) or (16), but the structure of the vacuum, along with the confinement properties, already pose important conceptual questions regarding initial conditions, Cauchy surfaces, and determinism. The main problem is that, practically and experimentally, we can claim to be able to fix initial conditions, at the fundamental level, in the perturbative, asymptotic states, that lie in ${\cal H}_{0}$, but only in composite confined states when it comes to ${\cal H}_{\mu}$, as explained above.
Obviously, since the entire evolution is unitary and Lorentz invariant (in fact, these are the basic requirements built in the algebraic construction described in the next Section) it remains so if the initial and final states are confined (protons, baryons, hadrons) but this is not checked at the level of quarks and gluons, only at the level of the composite initial and final states. Intermediate states and calculations also obey unitarity and causality, by (15) or (16), but the experimental setups, preparations and verifications are on composite states.
5 Generalizations
In order to gain more insight, trust the process, and venture in other demonstrations and generalizations of confining phenomena, it is useful to recall the axiomatic algebraic quantum field theory that uses operator algebras in the foundations of its construction.
[…]
A state of $\cal A$} is a linear functional, $\phi$, from $\cal A$} to the complex numbers, which is positive, $\phi(A^{*}A)\geq 0$, and normalized, $\phi(I)=1$, where $I$ is the unity element of the algebra.
[…]
Symmetries are expressed by elements, $g$, of a Lie group, $G$, and the action of a symmetry on the elements of the algebra transforms every operator $Q$ to $U_{g}QU_{g}^{*}$, with a unitary representation, $U_{g}$, of the group on the Hilbert space.
Quantum field theory considers the observables measured in the spacetime domain $D=V\times T$, of the finite three-dimensional volume, $V$, of a laboratory, over a finite period of time, $T$. These observables, ${\cal A}(D)$, over all compact subsets, $D$, of Minkowski spacetime, generate an abstract algebra satisfying the basic postulates of locality (on spacelike separated distinct subregions, they commute) and Poincare covariance.
The Poincare group, $\cal P$}, is the semidirect product,
${\cal P}={\cal T}\ltimes{\cal L},$ (17)
[…]
The translation group has a unitary representation $T(a)=e^{iP\cdot a}$, for any vector, $a^{\mu}$, of the Minkowski spacetime, and $P^{\mu}$ is the energy-momentum vector. Time evolution is done with the unitary Hamiltonian operator, $P^{0}$.
Operators that decrease the energy are defined as in Section 3, and the vacuum state, $\phi$, is defined, so that it satisfies $\phi(Q^{*}Q)=0$, for any operator, $Q$, which decreases the energy in some coordinate system.
Unitarity and Poincare invariance are built in the construction of the quantum field theory, which is generated with the GNS prescription from the vacuum state. In the case of the Yang-Mills theory, the vacuum state is not a pure vector state, it is mixed, and the resulting Hilbert space is a direct sum of Hilbert spaces built on the perturbative and the confining vacuum, $|\Omega_{0}>$ and $|\Omega_{\mu}>$, respectively.
In the confining vacuum we have (14)
$\Omega_{\mu}|T(Q_{1}(x_{1})Q_{2}(x_{2})...Q_{N}(x_{N}))|\Omega_{\mu}>=0,$ (18)
for any operators, at any spacetime points, $x_{i}$, unless they lie in a timelike worldline (with a spacelike width of the order of the scale of the theory). Spacelike confinement leads to a timelike evolution along, what is sometimes called, a worldline of an observer.
The notion of an observer is not well-defined in physics, it is sometimes used to denote a measuring device (a detector making measurements of physical observables) and sometimes to denote a conscious entity registering such events. I will argue, based on (14), that these two very different concepts use the property of confinement as a necessary (certainly not sufficient) condition.
6 Consciousness as a confined state
There are several competing theories of consciousness, and they sometimes have some common postulates. Instead of comparing and analysing various different approaches, I will first make some comments using some of the postulates of the integrated information theory of consciousness (IIT) [9] and my main point is to argue for confinement as a necessary condition, its relevance, and possibility of realization in a generalised context, not to present a new, or justify an existing theory of consciousness.
According to some of the postulates of IIT, consciousness has the properties of integration and exclusion. Namely, it is irreducible to separate elements, it has borders, and is unified. The integrated information of the system can be thought of as information that is not stored in spacelike separated partitions, but “integrated” in the system as a whole. It is also interesting to note other arguments of theories of consciousness that challenge views of memory as stored in spacelike places in the brain [10], and propose “temporal” structure and organization.
The above characteristics of some of the postulates of IIT and other theories of consciousness are naturally embodied in a confined state, with the basic property (14)
$\Omega_{\mu}|T(Q_{1}(x_{1})Q_{2}(x_{2})...Q_{N}(x_{N}))|\Omega_{\mu}>=0,$ (19)
unless the $x_{i}$ lie in a timelike worldline (with an appropriate spacelike width).
Neither IIT, nor any other theory of consciousness, can easily accomodate the cluster decomposition property that holds on asymptotic states built on a unique vacuum.
Interestingly, there have also been some arguments for theories of consciousness using quantum mechanical treatments [11]. These arguments use quantum mechanics, both conceptually and physiologically, in order to quantify and qualify information, but they do not consider confinement in the sense presented here.
I would like to extend the quantum arguments, using the more general case of the algebraic method presented in Section 5, with the dynamics of operator algebras and symmetries, such as (17).
Namely, operator algebras express interactions and dynamics that were considered to describe relativistic quantum mechanics, but this does not have to be their only realization.
One may entertain the possibility of more general dynamics (noncommutative and highly nontrivial, of course) and other associated symmetries (but with a “temporal” and “spatial” component) which are realized and admit a mixed vacuum state that enjoys spatial confinement.
The dynamics of the operator algebra, the symmetries ${\cal P}={\cal T}\ltimes{\cal L}$, their representation $T(a)=e^{iP\cdot a}$, and temporal evolution governed by $P^{0}$, do not have to correspond to a quantum interaction, or the Lorentz and Poincare group. They can be related to different interactions and symmetries (quantum, classical, neural, biological) as long as they have a generalised operator algebra realization, a mixed, confining vacuum state, and distinct temporal and spatial components.
Then, the state of spatial confinement, as expressed with relations similar to (14), after a period which would be the equivalent of the biological evolution, can be considered to lead to a state of temporal consciousness of the kind experienced.
Several conceptual, and also some metaphysical questions can be better understood with the paradigm of “asymptotic” and “confined” states suggested here. For example, one can claim that “robot” or “zombie” states (defined as states that are assembled from asymptotic data, and programmed to react in a multitude of external stimuli with the same response as a conscious state) are not conscious, since they are, by default, asymptotic, not confined. The initial value problems that were also mentioned at the end of Section 4 can also be related to conceptual and metaphysical questions regarding “free will”, since conscious entities use sensors for all communications, and these respond to asymptotic data. The confinement paradigm shows that practical setups and interactions involving only asymptotic states (photons and electrons, for example) cannot fully describe a confined state (neither its initial condition nor its evolution). Providing initial data on ${\cal H}_{0}$, does not fully determine the evolution on ${\cal H}_{0}\oplus{\cal H}_{\mu}$.
7 The measurement problem
Most interpretations and practitioners of quantum physics agree that the concept of quantum measurement is problematic, not fully explained, and perhaps not even described within the framework of the theory itself, possibly associated with a non-unitary evolution that cannot be accomodated by its premises.
Important steps towards resolving the problem of measurement and wavefunction collapse have been made with the concept of decoherence [12], which considers the evolution of the quantum system to be measured, with states $|\psi_{n}>$, together with the measuring device, with pointer states $|\chi_{n}>$, and a reservoir with states $|\phi_{n}>$.
After a period of interaction, the state
$\sum_{n}c_{n}|\psi_{n}>|\chi_{n}>|\phi_{n}>,$ (20)
evolves to the density matrix
$\sum_{n,n^{\prime}}c_{n}c^{*}_{n^{\prime}}<\phi_{n^{\prime}}(t)|\phi_{n}(t)>|\psi_{n},\chi_{n}><\psi_{n^{\prime}},\chi_{n^{\prime}}|,$ (21)
after a partial trace over the reservoir states. It is claimed that the large size of the reservoir causes the non-diagonal matrix elements $\phi_{n^{\prime}}(t)|\phi_{n}(t)>$ to decay to zero, and reduces the final state to a diagonal density matrix that corresponds to a statistical mixture.
The observation made here, based on (14), is that the desired relation, $\phi_{n^{\prime}}(t)|\phi_{n}(t)>=0$, for $n\neq n^{\prime}$, can be related to confinement. Although usually present, there is no need for a separate and “large” reservoir, these are the confined states of the different parts of the detector.
Here the different parts of the detector are considered confined by the usual strong interaction. There is no need for the generalised concepts of confinement described in the previous Section, and certainly no need for the concept of consciousness in order to make a measurement. Conscious entities, on the other hand, also have physical parts, consisting of ordinary matter and responding to external stimuli, and can certainly make measurements, which will also look like a statistical mixture to them.
One may look at this argument as a further justification of decoherence, or as a process that takes place along with decoherence, to the same effect. It can be related to the emergence of classical behavior and the collapse of the wavefunction, in a similar manner. One can also consider “metatheorems” such as the statement that we cannot build a detector out of just asymptotic matter, say, just electrons and photons; we also need the baryons that bind them.
8 Conclusion
As in the original work [1], the comments and results presented in [2, 3] and here are somewhat disparate. The reason is that confinement is related to a generalization of the basic axioms of quantum field theory, and this can be associated with several distinct phenomena.
It is also hoped that the problems of quantum gravity can be related to these investigations, that the constraints of the spin-$2$ gravitons can be treated in a similar manner as the spin-$1$, non-Abelian gauge bosons, and that strong gravitational and cosmological effects, involving black hole or de Sitter horizons, can be clarified.
The postulate of the uniqueness of the vacuum is made in the original formulation of axiomatic quantum field theory, in terms of correlation functions [4]. In the algebraic formulation [5, 6], briefly discussed here in Section 5, only the physical postulates of causality and Lorentz invariance are used in the foundations. The vacuum can be a mixed state, and the postulate of uniqueness is made in order to have the additional, equivalent property of cluster decomposition, which states that, when we pull a physical system apart, we reduce it to various, weakly interacting, spacelike separated components.
This is not what happens in the strong interaction. It is noted, however, that the strong interaction is part of the Standard Model, and many precise estimates have to be made, that involve the confining vacuum, combining contributions from theoretical, phenomenological, experimental, and lattice physics. The results of the work presented here can be thought of as justifications of older phenomenological models, such as the “bag model”, with the introduction of only a few parameters, the confining vacuum energy density and the mixing between the two vacua. The breakdown of conventional quantum field theory at the axiomatic level, starting from “axiom zero”, raises many problems and questions towards the treatment of the strong interaction, from the theoretical and axiomatic level of Yang-Mills theory, to the calculation of its effects at the various scales involved.
Note
There are no additional data and no conflicts of interest regarding this work.